Comet Halley is an irregularly-shaped comet which is visible from Earth every 75-76 years. For the purposes of this question, we will assume that the comet is spherical.

a) Takingthediameterofthecomettobe10.6km,findtheapproximate volume of the comet in cubic centimetres. You may use the fact that the volume of a sphere is 3

V = 43 π r where r is the radius.


1
Expert's answer
2021-05-17T17:34:30-0400
"V=\frac{4}{3}\pi r^3\\V=\frac{4}{3}\pi (5.3)^3=624\ km^3"